An ideal gas has the t P-4 atm, and 1-100°C. What is its final temperature # the pressure is reduced to 1 atm and the volume expands to 1000 cmh ec
Answer #1

No of moles are constant so by PV=nRT

PiVi/Ti = PfVf/Tf

T should be in kelvin ....so Ti = 100 + 273 = 373 K


4 * 525 / 373 = 1 X 1000/Tf

5.63 = 1000/Tf

Tf = 1000/5.63 =177.62 K

In degree C ...Tf = 177.62 - 273 = -95.38 degree C